60% of the people who live in the state make 50,000 a year. 50% of the people who live in the state own two cars. What's the percent of people who don't own two cars and make less than $50,000?
Answer #1

Let S shows the event that people who live in the state make 50,000 a year and let C shows event that people who live in the state own two cars. So we hve

P(S) = 0.60 and P(C) = 0.50

Assuming these two events are independent so

P(S and C) = P(S) P(C) = 0.60 *0.50 = 0.30

The percent of people who don't own two cars and make less than $50,000 is

P(S' and C') = 1 - P(S or C) = 1 - [ P(S) + P(C) - P(S and C) ] = 1 - [ 0.60 + 0.50 - 0.30] = 0.20

Answer: 20%